Explain how a bondholder who purchased a bond in a prior period is affected by a decrease in market rates. Address the bondholder’s situation with regard to income and bond value.
Answer #1

Bond coupon rate: The interest rate paid on face value on bond.

Market rate: The interest rate used to discount the cash flows of the bond based on timing.

Yield to maturity: The weighted average discount rate used to find out the price of the bond.

Example: On January 1st, 2019 , 5% coupon bond with face value 1000 INR, 3 years maturity is sold at par value. The market rate is 5%. Now after one year, on January 1st, 2020, the market rate is 6%, then the value of cash flows and bond value is shown as follows:

Year 1: price = 50/(1+.06)+50/(1+0.06)^2+1000/(1+0.06)^2

price = 981.666

if market rate of 5% is used instead of 6% of present market rate,

price = 50/(1+0.05)+50/(1+0.05)^2+1000/(1+0.05)^2

price = 1000

if market rate of 4% instead of 6% of present market rate,

price = 50/(1+0.04)+50/(1+0.04)^2+1000/(1+0.04)^2

price = 1018.86

From the above example, it is clear that when market rates decrease with respect to coupon rate, the value of the bond and cash flows increase from par value.